Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 154     Entry time: 24 Apr 2023, 02:45
Author: Ulrike Boesenberg 
Group:  
Subject: shift summary 

  • Beam came back quickly after cryo-incident with ok seeding performance (not very stable).
  • Finish 3381-Pudell beamtime with SRO/STO samples at time delays and laser fluences.
  • Decided not to scan the HXRSS energy but stay at current position
  • Mount samples for 3331-Robinson experiment and search for Bragg reflections, we could find the specular reflection for the two largest samples but not the magnetostrictive one.
  • The specular reflection for the 8x8um sample is currently aligned.

 

Problems:

  • AGIPD darks could not be pushed into the calibration pipeline, the problem was fixed quickly by DOC and OCD

ELOG V3.1.4-7c3fd00